Prove that the dot between two vectors is commutative not associative


1
Expert's answer
2021-06-14T15:11:39-0400

The dot product of two Euclidean vectors "\vec a"  and "\vec b"  is defined by


"\vec a \cdot \vec b=||\vec a||\ ||\vec b|| \cos \theta"

where "\theta"  is the angle between "\vec a" and "\vec b."


The dot between two vectors is commutative. This follows from the definition of the dot product


"\vec a \cdot \vec b=||\vec a||\ ||\vec b|| \cos \theta\ =||\vec b||\ ||\vec a|| \cos \theta=\vec b \cdot \vec a"

where "\theta" is the angle between "\vec a" and "\vec b."


The dot product between a scalar "(\vec a \cdot \vec b)" ) and a vector "(\vec c)" is not defined, which means that the expressions involved in the associative property, "(\vec a \cdot \vec b)\cdot \vec c"  or "\vec a \cdot (\vec b\cdot \vec c)" are both ill-defined.

Hence the dot between two vectors is not associative.